OpenStudy (loser66):

consider the RLC circuit with R = 2 (ohm) L = 1/2H; C = 2/5F . Initially the capacitor is uncharged, and no current is flowing in the circuit. Determine the current for t > 0 , if the applied EMF is E(t) = 50t if 0= pi

OpenStudy (loser66):

I know the differential equation for it is \[\frac{d^2q}{tt^2}+4\frac{dq}{dt}+5q=E(t)\] and solve for 2 cases of partial solution of E(t) but don't understand about the initial condition.

OpenStudy (loser66):

@Euler271

OpenStudy (loser66):

the right hand side should be E(t)/L

OpenStudy (anonymous):

initially uncharged : q(0) = 0 initially no curren: dq/dt(0) = 0 current is the rate of change of charge. so i = dq/dt

OpenStudy (loser66):

although i = dq/dt, we have to follow that format, right? I mean dq/dt to solve . if not, we get stuck at i' + i + q = ..., right?

OpenStudy (anonymous):

ya, its looked at as Li' + Ri + 1/Cintegral i dt = q you should be looking at it as a second order DE though. you could almost ignore what i said :P

OpenStudy (loser66):

XD

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/RLC_circuit if you want to see it :P

OpenStudy (loser66):

Thanks for killing me, heheheh.... (joking) I will read it carefully. thank you

OpenStudy (anonymous):

hehe :P you shouldn't really bother with it. not sure it's very solvable without software. i was using it in a "modelling and system analysis" class. was about setting up equations for real life models

OpenStudy (loser66):

I have quiz tomorrow and just this part to be "completely well prepare student" and it 's midnight here. hopefully I don't oversleep to wake up on time. hehehe

OpenStudy (anonymous):

lol you should set at least 2 alarms :P

OpenStudy (loser66):

ok, thank you. I am better go to bed now,

OpenStudy (anonymous):

good night and good luck ^_^

OpenStudy (loser66):

ty